The Department of Computer Science and Engineering

University of Nevada, Reno

cordially invites you to a Master’s colloquium

VFIRE: Virtual Fire in Realistic Environments A Framework for Wildfire Visualization in Immersive Environments

A thesis submitted in partial fulfillment of the requirements for the degree of Master of Science with a major in Computer Engineering.

by

Michael A. Penick

Abstract: Wildfire spread model output is used to make important and oftentimes expensive decisions. This thesis presents “VFire - Virtual Fire in Realistic Environment” an application and more importantly a framework for visualizing wildfire simulations in immersive environments. VFire will allow its users to visualize wildfires from perspectives and positions, which are normally too dangerous. Recent developments in graphics and virtual reality technology allow us to achieve this goal. It has recently become possible to visualize wildfire simulations in more realistic and immersive ways then has ever been achieved. To this end, VFire is an immersive visualization application that aids in wildfire training and data analysis endeavors and also a framework with which future wildfire applications can be developed.

3:00 pm, Wednesday, April 8, 2007

Scrugham Engineering and Mines (SEM) room 201

For more information contact Dr. Fred Harris @ 784-6571 (fredh@cse.unr.edu)